OpenStudy (31356):

The cone shown has a diameter of 18 meters and a slant height of 6 meters. Which choice is closest to the lateral surface area? Use 3.14 to approximate pi. MEDAL REWARDED!!

OpenStudy (31356):

A)170 B)424 C)339 D)1356

OpenStudy (jdoe0001):

\(\bf \textit{lateral area of a cone}=\pi r \cdot \textit{slant height}\\ \quad \\ r = radius = \frac{diameter}{2} \)

OpenStudy (31356):

Okay thanks!

OpenStudy (jdoe0001):

yw